Yellow Sand Banking has a weighted-average cost of capital of 7.08 percent and is evaluating two projects: A and B. Project A involves an initial investment of 6,250 dollars and an expected cash flow of 10,313 dollars in 4 years. Project A is considered more risky than an average-risk project at Yellow Sand Banking, such that the appropriate discount rate for it is 1.18 percentage points different than the discount rate used for an average-risk project at Yellow Sand Banking. The internal rate of return for project A is 13.34 percent. Project B involves an initial investment of 4,750 dollars and an expected cash flow of 8,835 dollars in 9 years. Project B is considered less risky than an average-risk project at Yellow Sand Banking, such that the appropriate discount rate for it is 1.44 percentage points different than the discount rate used for an average-risk project at Yellow Sand Banking. The internal rate of return for project B is 7.14 percent. What is X if X equals the NPV of project A plus the NPV of project B?
Answer #1

Given:

Weighted average cost of capital = 7.08%

Initial Investment of project A = $6,250

Expected cash flow of project A = $10,313 in 4 years

Discount rate = 1.18 points different than avg. risk project

Internal rate of return for Project A = 13.34%

Initial Investment of project B = $4,750

Expected cash flow of project A = $8,835 in 9 years

Discount rate = 1.44 points different than avg. risk project

Internal rate of return for Project A = 7.14%

NPV of project A = -6250 + 10313 / (1+7.08%+1.18%)^4

                             = -6250 + 10313 / (1.0826)^4

                             = -6250 + 10313 / 1.37

                             = -6250 + 7527.74

                             = 1277.74

NPV of project B = -4750 + 8835 / (1+7.08%-1.44%)^9

                             = -4750 + 8835 / 1.64

                             = 637.20

X = NPV of Project A + NPV of Project B

X = 1277.74 + 637.20

X = $1,914.94